The right circular cone has a volume = 24(pi=3.14)cubic cm and r = 3, What is the height?
Answer by malakumar_kos@yahoo.com(315)


volume = v = 24 cc
radius= r = 3.cms
height = h = ?
v = pi.r^2.h/3 = (3.14).3.3.h/3 = (3.14).3.h
24 = (3.14).3.h therefore h = 24/9.42 = 2.54cms
therefore h = 2.54cms